# LiDAR Remote Sensing Meets Weak Supervision: Concepts, Methods, and Perspectives

### v1

@@ 4 Perspectives and Future Work
LiDAR Remote Sensing Meets LM. The integration of LiDAR remote sensing with large-scale models offers significant potential for Earth observation. LiDAR provides high-resolution 3D spatial data, and when combined with pre-trained models for remote sensing images, it enables robust and scalable solutions for tasks such as biomass estimation and terrain analysis. Furthermore, integrating LiDAR data with multimodal large models (e.g., vision-language models) can address challenges such as parameter inversion and scene interpretation [239]. Cross-modal supervision allows the model to leverage text, imagery, and other information to provide supervisory signals for LiDAR tasks, enabling weak supervision at scale. For a given 3D pre-trained large-scale model in the future, it could be applied to Earth observation. Pre-training on diverse 3D Earth observations can significantly improve the model’s ability to address the inherent complexities and domain shift in remote sensing.
@@ 5 Conclusion
In summary, we present the first comprehensive overview of LiDAR remote sensing from a machine learning perspective. For LiDAR interpretation, we discuss key tasks such as semantic segmentation, instance segmentation, and object detection, emphasizing the latest methods and related weakly supervised learning techniques. In the context of LiDAR inversion, we focus on applications like canopy height estimation, building height estimation, and biomass inversion. Although weak supervision is increasingly being applied to inversion tasks, existing methods are still limited and have yet to fully harness the potential of weakly supervised learning.
We also propose several promising research directions for further exploration, including better utilization of widely available annotations, applying weak supervision to address domain shift in LiDAR remote sensing, developing new paradigms for LiDAR inversion, proposing new inversion datasets, and exploring large model development and generalized models for open-world environments. Through these explorations, the gap between point cloud interpretation and remote sensing applications cloud be bridged using machine learning (e.g., weakly supervised learning), advancing LiDAR remote sensing.
@@ Acknowledgements
This work was supported by the National Natural Science Foundation of China under Grant 42201481 and the National Key R&D Program of China (2024YFF0810400).

### v2

@@ 6 Conclusion
In summary, this paper presents the first comprehensive overview of LiDAR remote sensing from a weakly supervised learning perspective. We revisit two core directions—data interpretation and parameter inversion—under a unified machine learning framework. For LiDAR data interpretation, we review key tasks such as semantic segmentation and object detection, emphasizing representative WSL techniques. For LiDAR-based inversion, we examine applications such as canopy and building height estimation, where weak supervision is emerging but remains underexplored. These interpretation and inversion tasks face significant domain shift across regions, seasons, and sensors, highlighting the importance of domain generalization within WSL frameworks for scalable LiDAR remote sensing.
We further analyze LiDAR-specific challenges for WSL, such as irregular geometry and sampling sparsity, which demand tailored strategies beyond conventional 2D settings. Looking ahead, we outline several promising research directions: leveraging ubiquitous weak annotations, developing domain-adaptive WSL to mitigate domain shift, advancing inversion with new paradigms and datasets, and exploring foundation model integration to support open-world, scalable, and adaptive remote sensing. Through these explorations, we introduce an interdisciplinary perspective and conceptual framework that encourages the integration of machine learning, remote sensing, and geoscience for advancing LiDAR remote sensing.
@@ Acknowledgements
This work was supported by the National Natural Science Foundation of China under Grant 42201481 and 42271365.